OpenStudy (anonymous):

You run a small furniture business. You sign a deal with a customer to deliver up to 500 chairs, the exact number to be determined by the customer later. The price will be $110 per chair up to 400 chairs, and above 400, the price will be reduced by $0.25 per chair (on the whole order) for every additional chair over 400 ordered. a) What is the largest revenue your company can make under this deal? R=______ dollars b) What is the smallest revenue your company can make under this deal? R=_____dollars please explain! Thanks:)

hero (hero):

Piecewise Function: x = number of chairs ordered f(x) = 110x, if x < 400 = 110x - .25x if 400 < x < 500

OpenStudy (anonymous):

okay:) so from here, do you set them equal to each other? :/not too sure :(

hero (hero):

Finding maximum and minimum values of the function. You should start off with graphing the function given the ranges

OpenStudy (anonymous):

okay... so set it up on a x/y axis? ;/ not sure how to graph this though :(

hero (hero):

You graph each one at a time. You can use a calculator or desmos.com to graph. They are just linear functions anyway. Nothing difficult.

OpenStudy (anonymous):

i punched it into wolfram alpha... is this right? for 110x, if x < 400 ?

OpenStudy (anonymous):

but i couldn't get anything from that second function :/

hero (hero):

It is better to use desmos.com to graph because you can set up proper limits for x

hero (hero):

I'm surprised you made it this far without knowing how to graph on a calc.

OpenStudy (anonymous):

oh on my graphing calc?

OpenStudy (anonymous):

and don't know how to use desmos... never used it before :/

hero (hero):

Yes, you can do it on your graphing calc as well. Hang on a min...

OpenStudy (anonymous):

ahh okay:P yeah i just forget how to enter the domains :/

OpenStudy (anonymous):

ohhh okay i see now :) so what's the next step from here?

hero (hero):

You should be able to calculate the min and max amounts. You can assume that at least one chair will be sold for minimum. 500 chairs for maximum

hero (hero):

So at least $110 for minimum

hero (hero):

f(1) = 110

hero (hero):

Compute f(500) for maximum

OpenStudy (anonymous):

f(500)=110(500) =55,000 ?

hero (hero):

According to the piecewise function you use 110x - .25x if 400 < x < 500

hero (hero):

f(500) = 110(500) - .25(500)

OpenStudy (anonymous):

ohh so f(500) = 110(500) - .25(500) =55,000 - 125 = 54, 875 ?

hero (hero):

looks right

OpenStudy (anonymous):

and so would that be the largest revenue? since it's the max?

OpenStudy (anonymous):

and the smallest revenue is 110? so 54874=max and 110=min ?

hero (hero):

I should have used R(x) for the function

hero (hero):

Yes, that should be correct

OpenStudy (anonymous):

ahh okay!! yay! So this problem is all complete?! :O

hero (hero):

I would assume so unless there are other questions to answer

OpenStudy (anonymous):

no, that's all :) hehe yay! thanks!!

hero (hero):

The minimum revenue assumes the customer will purchase at least one chair.

OpenStudy (anonymous):

ahh okay:)

OpenStudy (anonymous):

thank you!! :D

hero (hero):

yw @iheartfood